# EmojiDiff: Advanced Facial Expression Control with High Identity Preservation in Portrait Generation

### v1

@@ Abstract
This paper aims to bring fine-grained expression control to identity-preserving portrait generation. Existing methods tend to synthesize portraits with either neutral or stereotypical expressions. Even when supplemented with control signals like facial landmarks, these models struggle to generate accurate and vivid expressions following user instructions. To solve this, we introduce EmojiDiff, an end-to-end solution to facilitate simultaneous dual control of fine expression and identity. Unlike the conventional methods using coarse control signals, our method directly accepts RGB expression images as input templates to provide extremely accurate and fine-grained expression control in the diffusion process. As its core, an innovative decoupled scheme is proposed to disentangle expression features in the expression template from other extraneous information, such as identity, skin, and style. On one hand, we introduce ID-irrelevant Data Iteration (IDI) to synthesize extremely high-quality cross-identity expression pairs for decoupled training, which is the crucial foundation to filter out identity information hidden in the expressions. On the other hand, we meticulously investigate network layer function and select expression-sensitive layers to inject reference expression features, effectively preventing style leakage from expression signals. To further improve identity fidelity, we propose a novel fine-tuning strategy named ID-enhanced Contrast Alignment (ICA), which eliminates the negative impact of expression control on original identity preservation. Experimental results demonstrate that our method remarkably outperforms counterparts, achieves precise expression control with highly maintained identity, and generalizes well to various diffusion models.
Figure 1: Expression customization by different methods. (a) The methods extract the control signals from expressions and then inject them into diffusion models. (b) The methods first generate expressionless images, and manipulate them through animation. (c) The proposed method directly incorporates portrait images and reference expressions into diffusion models, bypassing the limitations of the detection network and post-processing.
@@ 1 Introduction
At the same time, researchers seek to use reference images to manipulate portrait expressions during the diffusion process and have devoted many efforts [9, 23, 33]. The representatives of them, including MediaPipe [7], Arc2Face [32], and FineFace [42], focus on extracting control signals from expression images and then injecting them into the diffusion model. However, fine-grained expression control conflicts with identity preservation. Existing methods use coarse control signals (e.g., landmarks, poses) instead of original RGB expression images to prevent interference from identity-related information (such as skin and lightning). However, these signals lose many facial details (e.g. frowning, pouting, cheek twitching), which are also limited by the robustness and accuracy of third-party networks. Besides, the integration of the above methods with portrait generation approaches also affects the vanilla identity-preserving generation. This limitation arises because these methods overlook the relationship between expression control and identity control, resulting in facial muscle changes caused by expression control reducing identity fidelity.
@@
In this paper, we propose EmojiDiff, an innovative end-to-end solution that seamlessly integrates fine-grained expression control with high-fidelity portrait generation. To achieve accurate expression control, we argue that the model should derive the motion directly from the original reference expressions, permitting comprehensive expression representation and robust generalization [16, 45]. However, expression images themselves contain irrelevant information like identity and style. If the model directly replicates the appearance and structure from the expression reference, it compromises the identity and style of generated images (referred to as identity leakage and style leakage). To address these challenges, a decoupled scheme is proposed. On one hand, we innovate ID-irrelevant Data Iteration (IDI), a novel strategy for decoupled training. Unlike previous methods [2, 32, 42] that employ same-identity data pairs, IDI sequentially manipulates the expression and identity of the image to synthesize high-quality data pairs with consistent expressions but differing identities. Training the model with these cross-identity expression pairs facilitates the transfer of detailed expressions to the generated output while implicitly filtering out identity information concealed in the expressions. On the other hand, we discover that the network layers serve distinct functions, with specific layers responsible for incorporating expression features while omitting expression-irrelevant information. By injecting expression features solely into these layers, we effectively address most style leakage issues. To optimize the compatibility between expression and identity control, we develop the ID-enhanced Contrast Alignment (ICA) to further fine-tune the model. ICA efficiently computes the identity loss and expression loss to align identity representations of images generated with and without the expression control, thereby mitigating the negative impacts of expression control on identity and improving identity fidelity. The main contribution of this paper lies in four-fold:
1) We propose an end-to-end solution that integrates fine-grained expression control, high-fidelity ID preservation, and strong adaptability to various diffusion models.
2) We innovate a decoupled scheme to disentangle expression features from extraneous information in RGB expression images, which effectively prevents ID leakage and style leakage through ID-irrelevant Data Iteration and selective feature injection at expression-sensitive layers.
@@
3) We propose ID-enhanced Contrast Alignment, a fine-tuning strategy designed to ensure generated portraits maintain stable identity characteristics across diverse expressions, thereby improving identity fidelity.
@@
4) We extensively evaluate the performance of the proposed method on different base models and expressions, showing that EmojiDiff achieves better performance compared to state-of-the-art methods.
@@
Figure 2: Overview of the proposed method. To integrate expression control into diffusion models, we aim to train the model using cross-identity triplet data $\{\mathbf{S}_{k}^{g^{\prime}},\mathbf{R}_{k^{\prime}}^{g},\mathbf{D}_{k}^{g}\}$ and mitigate the negative impact on the original structure through contrastive alignment. The method involves four stages. First, the fundamental expression controller (i.e., Base E-Adapter) is trained with same-identity triplet data (the detailed structure of the E-Adapter is illustrated in Fig. 3). Next, the trained Base E-Adapter and FaceFusion [11] are utilized to alter the identity of portraits while maintaining consistent expressions, thereby creating cross-identity expression pairs $\{\mathbf{R}_{k^{\prime}}^{g},\mathbf{D}_{k}^{g}\}$. Subsequently, the Refined E-Adapter is trained using newly synthesized data, facilitating dual control of identity and expression without ID leakage. Finally, the Refined E-Adapter is fine-tuned by introducing expression and identity loss, further minimizing its negative impact on identity.
@@
Figure 3: The proposed E-Adapter. The embeddings of identity and expression images are obtained through their respective branches. Subsequently, the ID embedding and the text embedding are incorporated into all network layers through attention, as demonstrated in Eq. (3). In designated expression layers, all embeddings are integrated into the network, as depicted in Eq. (5).
@@ 2.2 Expression Customization
Most methods for expression animation integrate control signals extracted from expression images into diffusion models. ControlNet [49] is a pioneering effort that enables the incorporation of human poses, depth maps, and other image conditions. Following this framework, MediaPipe [7] specializes in extracting facial landmarks from reference expressions to preserve more expression details. Arc2Face [9] extract 3D normal maps [13] to personalize expressions based on given portraits. Additionally, some studies pursue more flexible control. Collaborative Diffusion [23] allows localized editing of facial parts using semantic maps from expression references. [33] combined valence & arousal, and action units and GANmut [8] to characterize emotion information. DiffSFSR [27] utilizes an emotional dictionary to transfer facial expressions to the generated image. FineFace [42] employs action units to allow localized control by describing the intensity of facial muscle movements.
In practice, portrait animation [48, 21, 30, 10, 4] serves as an indirect tool for expression customization. Notable examples, such as FaceAdapter [18], X-Portrait [45] and LivePortrait [16], transfer expressions from expression templates to portrait images. However, these methods cannot customize identities and expressions simultaneously, but instead post-process the images generated by the diffusion model, adversely affecting identity fidelity and image quality. Our method avoids the need for third-party networks and post-processing. Instead, it directly uses RGB expression images and portrait images to generate expressive images Fig. 1 illustrates the differences among these methods.
@@ 3 Methodology
Given a source portrait $\mathbf{S}$, a textual description $\mathbf{T}$, and a reference $\mathbf{R}$, our method aims to generate customized images $\mathbf{D}$. To enable the model to accept RGB expression inputs, we aim to construct a high-quality ternary data sets $\mathcal{T}=\{\mathbf{S}_{k}^{g^{\prime}},\mathbf{R}_{k^{\prime}}^{g},\mathbf{D}_{k}^{g}\}$, where $k$ and $g$ denote the $k$-th ID and the $g$-th expression, respectively. During training, the model takes $\mathbf{S}_{k}^{g^{\prime}},\mathbf{R}_{k^{\prime}}^{g}$ as inputs and $\mathbf{D}_{k}^{g}$ as the learning objective, represented as $\mathbf{D}_{k}^{g}=\mathcal{F}(\mathbf{S}_{k}^{g^{\prime}},\mathbf{R}_{k^{\prime}}^{g})$, where $\mathcal{F}$ is the trained expression controller. This setup allows the model to maintain the identity of $\mathbf{S}$ while accurately replicating the expression from $\mathbf{R}$. However, existing methods [45, 46] fail to synthesize high-quality cross-identity expression pairs  $\{\mathbf{R}_{k^{\prime}}^{g},\mathbf{D}_{k}^{g}\}$ in $\mathcal{T}$. Besides, the integration of expression control hinders the original ID control.
To overcome these issues, our method is divided into four stages: Firstly, we design the expression controller and train a base version (Base E-Adapter) using same-identity ternary data sets as detailed in Sec. 3.2. The images generated by Base E-Adapter exhibit a very high expression consistency with given expression references, although they also have a similar appearance to the references (ID leakage). Secondly, we develop the ID-irrelevant Data Iteration (IDI) to alter the identity of the images generated by Base E-Adapter, thus promoting the given reference expressions and the altered images form data pairs with the same expression but different identity, as outlined in Sec. 3.3. Using the newly created cross-identity dataset, we retrain the proposed expression controller to develop the Refined E-Adapter. Refined E-Adapter accurately controls expression details while preventing ID leakage when generating portrait images (see Sec. 3.4). Finally, we introduce ID-enhanced Contrast Alignment (ICA) to further fine-tune the Refined E-Adapter, reducing the negative impact of expression manipulation on ID control (Sec. 3.5). Fig. 2 provides an overview of our pipeline.
@@ 3.1 Preliminary
3.2 Base E-Adapter Training
Cross-identity expression pairs $\{\mathbf{R}_{k^{\prime}}^{g},\mathbf{D}_{k}^{g}\}$ are essential for training expression controllers, as they disentangle expression features in RGB expression images from other irrelevant information. Nonetheless, existing methods [16, 11, 45, 46] struggle to synthesize data pairs that maintain high expression consistency while exhibiting significant identity discrepancies, particularly with extreme expressions. Therefore, we seek to develop a novel strategy for data construction.  Reviewing previous works [45, 46], we notice that they avoid using same-identity expression pairs $\{\mathbf{R}_{k}^{g},\mathbf{D}_{k}^{g}\}$ for controller training to prevent identity leakage during inference.  However, our insights suggest that the controller trained with same-identity pairs exhibits superior expression transfer capabilities, and it is highly suitable for data construction when combined with portrait generation methods. Specifically, we divide the synthesis of cross-identity expression pairs into two sequential steps. Initially, we utilize same-identity data to train an expression controller that accurately transfers reference expressions onto generated images without considering identity leakage. Subsequently, we apply portrait generation techniques to further alter the identity of the controller-generated image. With the powerful generative abilities of the diffusion model, the modified image forms a high-quality cross-identity data pair with the given expression template.
In this section, we first propose the expression controller (E-Adapter) suitable for portrait generation and then train its basic version for subsequent data construction. Inspired by IP-Adapter [47], the designed E-Adapter integrates identity and expression through an attention mechanism, ensuring the controller remains lightweight. As illustrated in Fig. 3, E-Adapter includes two main control branches: the identity and expression branches, which process portrait and expression images, respectively. In the identity branch, we inherited IP-Adapter to encode portrait images into identity embeddings and inject them into the U-Net along with text embeddings using a decoupled attention, formulated as:
@@ 3.2 Base E-Adapter Training
In the expression branch, the local image is preprocessed from the origin expression reference, which removes irrelevant elements such as the background, face contour, etc. We then extract the expression embedding from the local image and inject it into the same cross-attention module with the identity embedding, denoted as:
@@
Figure 4: Visualization of the attention maps, where the regions colored red/blue represent the most/least salient areas.
Noteworthy, we discover that expression images contain some abstract, expression-irrelevant information (e.g., lighting, skin), which may affect the quality and style of the generated image. As shown in Fig. 4, we visualize different cross-attention layers in the adapter. It becomes evident that the cross-attention layers serve distinct responsibilities, some of which are responsible for injecting facial information and showing pronounced responses in the facial region. Furthermore, certain layers are sensitive to structural information while neglecting style information, making them ideal for expression control. Consequently, we inject expression embeddings exclusively into these layers (i.e., Expression Layer), effectively avoiding style leakage.
@@
3.3 ID-irrelevant Data Iteration
The image identity constructed directly using the Base E-Adapter is greatly influenced by the expression reference’s appearance. The core idea of IDI is to minimize this disruption to further increase the identity distinction between the generated image and the expression reference. Initially, we innovatively reuse the ID control branch of the base E-Adapter to modify the identity of the generated image. Specifically, we utilize expression images $\mathbf{R}_{k}^{g}$ inherited from training as the expression condition but randomly selected portraits as the input of the identity branch, combing with the diffusion model to synthesize the restruct image. Since the Base E-Adapter has been exposed to all expression images, it effectively transmits expressions to the restruct images. Through the identity branch, we also integrate the identity of randomized portraits. Notably, the identity of the generated image does not need to precisely match the given ID condition; it simply needs to differ from the identity of the expression image. Additionally, we employ FaceFusion [11] on the restruct image to create the faceswap image, further increasing the identity disparity between the generated portrait and the expression reference. Occasionally, the whole process might modify expressions (e.g., changing from blinking to non-blinking), and we employ facial blendshapes [15] and landmark [29] differences to exclude expression-changed data.
Through these processes, we acquire cross-identity expression pairs $\{\mathbf{R}_{k^{\prime}}^{g},\mathbf{D}_{k}^{g}\}$ that maintain highly consistent expressions while exhibiting completely distinct identity, and the example is illustrated in Fig. 5.
Figure 5: Examples of Stage II.
3.4 Refined E-Adapter Training
Utilizing the newly developed cross-identity data, we train a new expression controller termed the Refined E-Adapter. The key distinction from the Base E-Adapter training is that the expression and identity are no longer sourced from the same portrait. Consequently, the identity and expression control branches are implicitly decoupled in their functions, each responsible solely for conveying identity and expression, respectively.

### v2

@@ Abstract
This paper aims to bring fine-grained expression control while maintaining high-fidelity identity in portrait generation. This is challenging due to the mutual interference between expression and identity. On one hand, fine expression control signals inevitably introduce appearance-related semantics (e.g., facial contours, and ratio), which impact the identity of the generated portrait. On the other hand, even coarse-grained expression control can cause facial changes that compromise identity, since they all act on the face. These limitations remain unaddressed by previous generation methods, which primarily rely on coarse control signals or two-stage inference that integrates portrait animation. Here, we introduce EmojiDiff, the first end-to-end solution that enables simultaneous control of extremely detailed expression (RGB-level) and high-fidelity identity in portrait generation. To address the above challenges, EmojiDiff adopts a two-stage scheme involving decoupled training and fine-tuning. For decoupled training, we innovate ID-irrelevant Data Iteration (IDI) to synthesize cross-identity expression pairs by dividing and optimizing the processes of maintaining expression and altering identity, thereby ensuring stable and high-quality data generation. Training the model with this data, we effectively disentangle fine expression features in the expression template from other extraneous information (e.g., identity, skin). Subsequently, we present ID-enhanced Contrast Alignment (ICA) for further fine-tuning. ICA achieves rapid reconstruction and joint supervision of identity and expression information, thus aligning identity representations of images with and without expression control. Experimental results demonstrate that our method remarkably outperforms counterparts, achieves precise expression control with highly maintained identity, and generalizes well to various diffusion models. Code and data will be publicly available at https://emojidiff.github.io/.
Figure 1: Expression customization by different methods. (a) The methods extract the control signals from expressions and inject them into diffusion models. (b) The methods generate stylized images and manipulate them through animation. (c) The proposed method simultaneously incorporates portrait images and reference expressions, generating stylistic images in an end-to-end manner.
@@ 1 Introduction
Researchers [10, 24, 34] seek to use reference images to manipulate portrait expressions. However, due to the strong coupling between expression and identity, it is challenging to introduce precise expression control while preserving high identity fidelity. First, fine-grained expression control conflicts with identity preservation. Excessively detailed expression templates also bring appearance-related semantic and style information (such as facial contours and skin). If the model directly replicates the appearance and structure from the expression reference, it compromises the identity of generated images (referred to as ID leakage). Second, a potential conflict exists between facial expression and identity control since both manipulate the face. Expression control can modify the facial topology and reduce identity fidelity, while identity control may make expressions toward neutrality. Third, although the highly flexible diffusion model and Style LoRA enrich the generated portraits, they also bring challenges in the generalizability of identity and expression control across various stylized scenes.
Pioneering works can hardly deal with these challenges altogether. As depicted in Fig. 1, signal-conditioning generation methods [10, 24, 34] tend to extract and inject coarse expression signals (e.g., landmarks, poses) into the diffusion model to prevent ID leakage. However, these signals lose many facial details (e.g. frowning, pouting, and cheek twitching), and they are significantly limited by the accuracy of third-party networks, such as pose [2] and action unit detectors [3]. Portrait animation [22, 5] can serve as a post-processing tool for expression customization, which transfers expressions from reference templates to portrait images. However, these methods heavily rely on generated stylized images, where the identity and data distribution of the portraits significantly deviate from the real world, negatively impacting identity fidelity, expression control, and image quality.
@@
In this paper, we introduce EmojiDiff, an innovative end-to-end solution that integrates fine-grained expression control with high-fidelity portrait generation. Unlike previous signal conditioning-based methods [2, 33, 43], we still insist that the model should derive the motion directly from the original RGB expressions, ensuring comprehensive expression representation and robust generalization [17, 47]. To achieve this, we propose a pluggable expression controller (E-Adapter) alongside its decoupled training strategy. As the core of the training, we innovate ID-irrelevant Data Iteration (IDI) to synthesize the identity-expression decoupling data (i.e., cross-identity expression pairs), which are rarely available in real-world data. IDI creatively divides cross-identity expression transfer into two steps: maintaining expressions and modifying identities. By sequentially optimizing the two steps, we achieve a stable and high-quality data supply. Training the controller with these pairs facilitates the transfer of detailed expressions to the generated output while implicitly filtering out identity information concealed in the expressions, thereby preventing ID leakage. To address the potential conflict between expression and identity, we develop the ID-enhanced Contrast Alignment (ICA) for fine-tuning. ICA efficiently computes the identity and expression loss by one-step reconstruction, aligning identity representations of images across various expressions. This process mitigates the negative impacts of expression control on identity and enhances identity fidelity. In summary, our contributions are fourfold:
1) To the best of our knowledge, EmojiDiff is the first end-to-end framework that achieves both fine-grained expression control and high ID fidelity portrait generation, accommodating various diffusion models and style LoRA.
2) We innovate ID-irrelevant Data Iteration for decoupled training, effectively disentangling fine expression features from RGB expression images and preventing ID leakage effectively.
@@
3) We propose ID-enhanced Contrast Alignment, an efficiently fine-tuning strategy designed to ensure generated portraits maintain stable identity characteristics across diverse expressions.
@@
4) We extensively evaluate the performance of the proposed method on different base models, styles, and expressions, showing that EmojiDiff achieves better performance compared to state-of-the-art methods.
@@
Figure 2: Overview of the proposed method. To integrate RGB-driven expression control into diffusion models, we aim to synthesize cross-identity data $\{\mathbf{S}_{k}^{g^{\prime}},\mathbf{R}_{k^{\prime}}^{g},\mathbf{D}_{k}^{g}\}$ for the model’s decoupled training, and mitigate the negative impact on the original identity through contrastive alignment fine-tuning. Before decoupled training, the fundamental expression controller (i.e., Base E-Adapter) is trained with same-identity triplet data to obtain expression transfer capabilities (the detailed structure of the E-Adapter is illustrated in Fig. 3). Next, the trained Base E-Adapter and FaceFusion [12] are utilized to alter the identity of portraits, thereby creating cross-identity expression pairs $\{\mathbf{R}_{k^{\prime}}^{g},\mathbf{D}_{k}^{g}\}$. Subsequently, the Refined E-Adapter uses newly synthesized data for disentangled training, facilitating dual control of identity and expression without ID leakage. Finally, the Refined E-Adapter is fine-tuned by expression and identity loss based on ANI.
@@
Figure 3: The proposed E-Adapter. The embeddings of identity and expression images are obtained through respective branches. Subsequently, the ID embedding, text embedding, and expression embeddings are integrated into the network, as depicted in Eq. (5).
@@ 2.2 Expression Customization
Most methods for expression animation integrate control signals extracted from expression images into diffusion models. ControlNet [51] is a pioneering effort that enables the incorporation of human poses, depth maps, and other image conditions. Following this framework, MediaPipe [8] specializes in extracting facial landmarks from reference expressions to preserve more expression details. Arc2Face [10] extracts 3D normal maps [14] to personalize expressions based on given portraits. Additionally, some studies pursue more flexible control. Collaborative Diffusion [24] allows localized editing of facial parts using semantic maps. [34] combined valence & arousal, and action units and GANmut [9] to characterize emotion information. DiffSFSR [28] utilizes an emotional dictionary to transfer facial expressions to the generated image. FineFace [43] employs action units to allow localized control by describing the intensity of facial muscle movements.
In practice, portrait animation [50, 22, 31, 11, 5, 45, 46] can adjust portrait expressions without altering the image background. Notable examples, such as FaceAdapter [19], X-Portrait [47], and LivePortrait [17], which can be used for post-processing of generated portrait images. However, these methods cannot simultaneously customize real-world identities and expressions, adversely affecting identity fidelity and image quality.
@@ 3 Methodology
Given a source portrait $\mathbf{S}$, a textual description $\mathbf{T}$, and a reference $\mathbf{R}$, our method aims to generate customized images $\mathbf{D}$. To enable the model to accept RGB expression inputs, we aim to construct high-quality ternary data sets $\mathcal{T}=\{\mathbf{S}_{k}^{g^{\prime}},\mathbf{R}_{k^{\prime}}^{g},\mathbf{D}_{k}^{g}\}$ for decoupled training, where $k$ and $g$ denote the $k$-th ID and the $g$-th expression, respectively. During training, the model takes $\mathbf{S}_{k}^{g^{\prime}},\mathbf{R}_{k^{\prime}}^{g}$ as inputs and $\mathbf{D}_{k}^{g}$ as the learning objective, represented as $\mathbf{D}_{k}^{g}=\mathcal{F}(\mathbf{S}_{k}^{g^{\prime}},\mathbf{R}_{k^{\prime}}^{g})$, where $\mathcal{F}$ is the trained expression controller. This setup allows the model to maintain the identity of $\mathbf{S}$ while accurately replicating the expression from $\mathbf{R}$. However, we are unable to collect qualified cross-identity expression pairs $\{\mathbf{R}_{k^{\prime}}^{g},\mathbf{D}_{k}^{g}\}$ from real-world scenarios. Besides, the integration of expression control hinders the original ID control.
To overcome these issues, our method is divided into two stages: decoupling training and fine-tuning. Before decoupled training, we construct the required data smartly in two steps. Firstly, we design the expression controller and train a base version (Base E-Adapter) using same-identity ternary data sets as detailed in Sec. 3.2.1. The images generated by Base E-Adapter exhibit a very high expression consistency with given expression references, although they also have a similar appearance to the references (ID leakage). Secondly, we develop the ID-irrelevant Data Iteration (IDI) to alter the identity of the images generated by Base E-Adapter, thus promoting the given reference expressions and the altered images to form data pairs with the same expression but different identity (as outlined in Sec. 3.2.2). Using the newly created cross-identity dataset, we perform decoupled training on the proposed expression controller to develop the Refined E-Adapter. Refined E-Adapter accurately controls expression details while preventing ID leakage when generating portrait images (see Sec. 3.2.3). Finally, we introduce ID-enhanced Contrast Alignment (ICA) to further fine-tune the Refined E-Adapter, reducing the negative impact of expression manipulation on ID control (Sec. 3.3). Fig. 2 provides an overview of our pipeline.
@@ 3.1 Preliminary
3.2 Decoupled Training
Cross-identity expression pairs $\{\mathbf{R}_{k^{\prime}}^{g},\mathbf{D}_{k}^{g}\}$ are essential for training expression controllers. However, we cannot obtain adequate samples from the real world due to the subtle differences in individual expressions. Additionally, existing methods [17, 12, 47, 48] also struggle to synthesize qualified data pairs, particularly with extreme expressions. Therefore, we seek to develop a novel strategy for data construction.  Reviewing previous works [47, 48], we notice that they avoid using same-identity expression pairs $\{\mathbf{R}_{k}^{g},\mathbf{D}_{k}^{g}\}$ for controller training to prevent identity leakage during inference. However, our insights suggest that the controller trained with same-identity pairs exhibits superior expression transfer capabilities, as shown in Fig. 4. Based upon this, we divide the synthesis of cross-identity expression pairs into two sequential steps: maintaining expression and modifying identity. Initially, we utilize same-identity data to train an expression controller that accurately transfers reference expressions into generated images without considering identity leakage. Subsequently, we apply portrait generation techniques to further alter the identity of the controller-generated image. With the powerful abilities of the diffusion model, the modified image and the given expression template form a high-quality cross-identity data pair. After that, the data can be used to train the model to decouple expressions from identities.
3.2.1 Base E-Adapter Training
In this subsection, we first propose the expression controller (E-Adapter) suitable for portrait generation and then train its basic version for subsequent data synthesis. Given the strong coupling between identity and expression, we employ a decoupled parallel attention to integrate them and design the E-Adapter, building upon IP-Adapter [49]. As illustrated in Fig. 3, E-Adapter includes two main control branches: the identity and expression branch. This simple yet effective design achieves three objectives. First, the parallel control of expression and identity reduces operation errors compared to the serial way. Second, the model can efficiently retrieve features from identity or expression, and learn how to incorporate the expression into the given identity. Third, the training module fully decouples expression and identity control, enabling independent or simultaneous manipulation of both (more empirical analysis can be found in Sec. 6). Specifically, in the identity branch, we inherited IP-Adapter to encode identity embeddings and inject them into the U-Net along with text embeddings, formulated as:
@@ 3.2.1 Base E-Adapter Training
In the expression branch, we first preprocess the original expression reference to obtain a local image, which removes irrelevant elements such as the background, face contour, etc. We then extract the expression embedding and inject it into the same cross-attention module with the identity embedding, denoted as:
@@
3.2.2 ID-irrelevant Data Iteration
The image identity constructed directly using the Base E-Adapter is greatly influenced by the expression reference’s appearance. The core idea of IDI is to minimize this disruption. To achieve this, we innovatively reuse Base E-Adapter to modify the identity of the generated image while keeping the expression unchanged. Specifically, we utilize expression images $\mathbf{R}_{k}^{g}$ inherited from training as the expression input but randomly selected portraits as the input of the identity branch, combining with the diffusion model to synthesize the restruct image. Subsequently, we perform FaceFusion [12] on the restruct image to create the faceswap image, further increasing the identity disparity between the generated image and the expression reference. Meanwhile, we employ facial blendshapes [16] and landmark [30] differences to exclude unqualified data (detailed in supplementary materia). Through these processes, we acquire cross-identity expression pairs $\{\mathbf{R}_{k^{\prime}}^{g},\mathbf{D}_{k}^{g}\}$ that maintain highly consistent expressions while exhibiting completely distinct identity, and the example is illustrated in Fig. 4.
Why can IDI keep the expression unchanged while changing the image’s identity? On one hand, as our insights at the beginning of Sec. 3.2, E-Adapter trained with the same identity data exhibits superior expression transfer capability, since the model tends to replicate the appearance and texture of the reference image. Additionally, the expression branch of the Base E-Adapter has been overfitted to all expression images by training. Consequently, IDI can effortlessly transmit the expression into the restruct images. On the other hand, the ID branch possesses certain identity control capabilities. The key is that the identity of the generated image does not need to match the input ID condition precisely; it simply needs to differ from that of the expression condition. By reusing the ID branch, we inject high-level semantic features of randomized portraits (e.g., gender, hairstyle, and face shape) into the restruct images. Moreover, the FaceFusion operation further alters the low-level facial textures, ensuring that the identity of the faceswap image is entirely distinct from the input.
Figure 4: Examples of IDI, detailed in Sec. 3.2.2.
3.2.3 Refined E-Adapter Decoupled Training
Utilizing the newly developed cross-identity data, we train a new expression controller termed the Refined E-Adapter. The controller’s expression and ID branches accept input from different portraits respectively. Consequently, the identity and expression control branches are explicitly decoupled in their functions, each responsible solely for conveying identity and expression, respectively.